# Contrasting exchange-field and spin-transfer torque driving mechanisms in all-electric electron spin resonance

### v1

@@ II.2 Quantum master equation
$\Gamma_{vl,ju,\alpha}^{-}(t)$ $=$ $\sum_{\sigma}\lambda_{lv,\sigma}^{*}\lambda_{ju,\sigma}\gamma_{\alpha\sigma}(t){\cal I}_{\alpha}^{-}(\Delta_{ju}),$ (8)
where the Fermi distribution function $f_{\alpha}(\epsilon)=1/\left(e^{\beta_{\alpha}(\epsilon-\mu_{\alpha})}+1\right)$, being $\beta_{\alpha}=1/k_{B}T_{\alpha}$ and $\mu_{\alpha}$ the chemical potential, is encoded in the integrals ${\cal I}^{\pm}_{\alpha}(\Delta_{ju})$ which are of the form:
[…]
@@
Notice that if $\hbar\omega\ll|\Delta_{ju}|$, Eq. 9 simplifies to [^2] A feasible assumption since $\hbar\omega$ will be in the 0.1 meV range or below, while $\Delta_{ju}\sim 100$ meV. This is because, in the rates only, $\Delta_{ju}$ always represents an energy difference between charged states with different numbers of electrons; otherwise, the rate is zero due to the definition of $\lambda_{lj,\sigma}$. Conversely, $\Delta_{lj}$ in Eq. 5 can be in the $\hbar\omega$ range.:
[…]
approximation valid for all the simulations in this work. Equation 10 contains $f^{+}_{\alpha}(\epsilon)=f_{\alpha}(\epsilon)$ and $f^{-}_{\alpha}(\epsilon)=1-f_{\alpha}(\epsilon)$ following [38]. Additionally, in Eqs. 7 and 8, we have introduced $\gamma_{\alpha\sigma}(t)=\left(1+A_{\alpha}\cos(\omega t)\right)^{2}\gamma_{\alpha\sigma}^{0}$ where $\gamma_{\alpha\sigma}^{0}=2\pi|t^{0}_{\alpha}|^{2}D_{\alpha}(1/2+\sigma P_{\alpha})=\gamma_{\alpha}^{0}(1/2+\sigma P_{\alpha})$ represents the level broadening due to the time-independent hopping, $t^{0}_{\alpha}$, for spin $\sigma$ in electrode $\alpha$. Further details on the transition rates, the derivation of the quantum master equation (QME), and its solution in the long-time limit can be found in [35].
It is important to mention that $\hbar/\tau_{c}$, up to first order in perturbation theory, should be infinitesimal in Eqs. 9 and 10. However, in this work, $\hbar/\tau_{c}\ll|\Delta_{ju}|\approx|\varepsilon|\ \mathrm{or}\ \varepsilon+U$ but nonzero, allowing us to explore two distinct regimes for the sequential tunneling: the one above the thresholds $|\varepsilon|,\varepsilon+U$ where an full electron can hop in or out from the QI and the one below the thresholds where the QI is in the so called Coulomb blockade (CB) regime and we tunnel through the tail of the resonance [36]. Therefore, the transition between both regimes is controlled by the DC voltages and happens when it equalizes the ionization energy $\varepsilon$ or charging one $\varepsilon+U$, which correspond to the situation of removing and adding one electron on the QI respectively.
We are aware of the limitations a finite $\hbar/\tau_{c}$ introduces, as a proper characterization of the Coulomb blockade (CB) requires the use of cotunneling transport or a Kondo scattering Hamiltonian [39, 27]. However, as long as $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$, we can overcome this limitation of the model and explore, to some extent, the implications of the FLT regime that governs the ESR signal in the CB regime. In this context, we expect low current at low DC biases, consistent with previous studies [38, 62, 39, 63, 42]. Future extensions of this model will aim to circumvent this issue for a better characterization of the system [^3] Going beyond a linear QME is likely required for this step, making the mathematical treatment of the model more challenging. However, first-order perturbation theory already provides rich results that follow experimental evidence, as we will explore in the next sections.
@@ II.3 Expression for the current
The expression for the current contains both DC and AC terms as a consequence of the hopping modulation. The measurable quantity in the experiment is the electric current signal in the long-time limit, which depends on the resonance condition as well as the relative orientation of the QI spin to the direction of the tip spin polarization. We refer to this current as ESR signal in the following. The coherences (off-diagonal elements of the density matrix) are explicitly included in the expression and are crucial for the correct calculation of the ESR signal, as pointed out in a previous work [36]. The coherences in Eq. 11 also lead to the so called homodyne detection of the electric current, a product of the misalignment of the tip’s spin and the spin of the impurity under it [34, 21, 65]. If the populations of the transient states ($0,2$) are negligible, we arrive at the following expression for the current:
$I_{L}(t)$ $=$ $\frac{e}{\hbar}\gamma_{L}\left(1+A_{L}\cos(\omega t)\right)^{2}$
$\times$ $\left[(1-2P_{L}\langle S_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{2g})\right\}\right.$
-- $\left.(1+2P_{L}\langle S_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{0g})\right\}\right]=-I_{R}(t).$
@@
$I_{L}=-I_{R}$ is due to electric current conservation, more details on the current can be found in Appendix C. In Eq. 12, we have employed the notation $g$ to refer to the ground state for a more general treatment instead of $\downarrow$, accounting for the situation in which the spin is tilted by the magnetic field with respect to the spin polarization on $z$. We do this because the states are always defined relative to the spin’s quantization axis, see Appendix A. A consequence of choosing the $z$-axis as the magnetization direction of the leads, where the spin polarization lies, the electric current becomes a measurement of the expectation value $\langle S_{z}\rangle$ as Eq. 12 shows. In general, for any magnetic orientation of the spin polarized lead, we expect that the current will measure the spin component parallel to it.
@@
Assuming that the magnetic field lies in the $xz$ plane, $\mathbf{B}=B(\sin\theta,0,\cos\theta)$, we have:
$\langle S_{z}\rangle$ $=$ $\frac{\hbar}{2}(\rho_{ee}-\rho_{gg})\cos\theta+\frac{\hbar}{2}(\rho_{ge}+\rho_{eg})\sin\theta.$ (13)
@@
$\theta$, also referred to as homodyne angle, is the angle formed between the direction of the left electrode’s magnetization, fixed along the $z$-axis, and the impurity spin when the $g$-factor is isotropic. We should emphasize that in Eq. 13, the main contribution to the populations comes from the Floquet number $n=0$, while the contribution from the coherences arises from $n=1$ [36]. Thus, the coherences are intrinsically time-dependent in the long-time limit. For the sake of completeness, the in-plane components of the spin expectation value are:
@@
The reader might wonder why the expectation value of the spin along $z$ depends on the coherences and not only on the difference in populations or why the $x$-component depends on the population. The reason is the inclusion of mixed states as a consequence of the magnetic field being tilted from the $z$ axis. Mathematically, Eq. 13 arises from changing the basis set from the classical up and down states of the Pauli matrix to the eigenstate basis set of the quantum impurity (QI) since ${\hat{S}}_{i}$ has to be expressed in the same basis as the reduced density matrix, meaning that we cannot simply use the standard Pauli matrices representation. This can be interpreted as a projection of the populations and coherences defined in the spin’s quantization axis onto the Cartesian coordinates. Equivalently, one could rotate the cartesian coordinates to the magnetic field direction, chaging the basis of the reduced density matrix instead, a procedure done in the literature [34, 21, 24, 18, 66] We emphasize that $\langle\mathbf{S}\rangle$ describes a quantum-statistical average and every component can adopt any number between -1/2 and 1/2.
Introducing Eq. 13 into Eq. 12 allows us to obtain an electric current equation that clearly shows the dependence on $\theta$. To simplify the result, we place ourselves in the long time where the populations are constant, i.e. $\rho_{ee(gg)}=P_{e(g)}$. Additionally, we consider that the linear driving term dominates ($A_{L}^{2}\sim 0$) so that the coherences are continuously Larmor precessing only, i.e. $\rho_{ge}(t)=\rho_{ge,-1}e^{i\omega t}+\rho_{ge,1}e^{-i\omega t}$, where the ±$\pm 1$ in the coherences refers to its Floquet number [36]. This yields:
@@
-- $P_{L}\left[(P_{e}-P_{g})\cos\theta-A_{L}(\rho_{ge,1}+\rho_{eg,-1})\sin\theta\ \right]$
@@
$I_{L}$$\approx$ $I_{0}-P_{L}\left[\Delta P_{eg}\cos\theta\right.\left.-A_{L}(\rho_{ge,1}+\rho_{eg,-1})\sin\theta\ \right]I_{0}^{\prime},$
where the new definitions $I_{0},\ I_{0}^{\prime}$, which contain a non-trivial DC bias dependence, see Eqs. 9 and 10, and $\Delta P_{eg}$ can be deduced by comparing terms.
Equation LABEL:homo_current_final shows the clear importance of the spin polarization: if $P_{L}=0$, the electric current will have a constant unpolarized value due to electrons or holes hopping from the left to right (or vice versa) passing through the QI. However, this electric current does not depend on the populations and coherences of the QI, making the current measurement incapable of detecting changes in the QI state and, by extension, it cannot cause any spin flip or resonant behavior even if there is a driving applied. This happens because only a non-zero spin polarization imbues an imbalance in the populations for the up and down states and a change in the tunneling magnetoresistance. If $P_{L}\neq 0$, a polarized current emerges and $\langle S_{z}\rangle$ starts to play a role. On resonance and in the long time limit, the population of both singly occupied levels will be $1/2$, assuming finite but small coherences. In such a situation, $\langle S_{z}\rangle\rightarrow 0$. On the other hand, if $P_{L}\neq 0$ but $A_{L}=0$, the QI will have a net spin polarization $\langle S_{z}\rangle\neq 0$. However, the states cannot exchange population since the off coherences are zero, leading to no resonance at the Larmor frequency. Another key insight from Eq. LABEL:homo_current_final is the dependence on the homodyne angle. Just as in the absence of driving, when $\theta=0$, no exchange occurs between the populations since $\sin\theta=0$. In contrast, at $\theta=\pi/2$, the electric current becomes independent of the populations, resulting in an ESR signal governed purely by the coherences. Thus, Eq. LABEL:homo_current_final clearly demonstrates that homodyne detection arises from both the misalignment between the tip and impurity spin and the presence of AC driving.
@@ II.4 Mapping onto a spin problem
Following Ref. [38], we now aim to map the QME (Eq. 5) onto a system of equations for the spin. This will become particularly useful for building some physical intuition on the definitions of the Rabi rate and the renormalization of the resonance energy. Using Eqs. 1314 and 15, the system of equations for the spin components and the population of the transient states is [^4] Equations LABEL:spin_equations assume a negligible Zeeman energy in the rates compared to the ionization and charge energy, $\varepsilon$ and $\varepsilon+U$. Since these energy thresholds are $\sim 100$ meV or more while $\Delta_{eg}=g\mu_{B}B<0.1$ meV, this is valid in the STM-ESR context  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25]. This allows us to write $\sum_{\alpha}P_{\alpha}\left[\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})\right]\sin\theta=\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))$, which further simplifies the spin system of equations; see Appendix D.:
$\hbar\langle\dot{S}_{z}\rangle$ $=\Delta_{ge}\langle{S}_{y}\rangle\sin\theta-\hbar\langle{S}_{z}\rangle/\tau_{\mathrm{rel}}+\hbar\langle\dot{S}_{z}\rangle_{\mathrm{acc}}$ (17)
$\hbar\langle\dot{S}_{x}\rangle$ $=\left[2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta-\Delta_{ge}\cos\theta\right]\langle{S}_{y}\rangle-\hbar\langle{S}_{x}\rangle/\tau_{\mathrm{rel}}$
$\hbar\langle\dot{S}_{y}\rangle$ $=-2\mbox{Im}(\Gamma_{ev,vg}(t))\langle S_{x}\rangle/\sin\theta-\hbar\langle{S}_{y}\rangle/\tau_{\mathrm{rel}}-\Delta_{ge}\left(\langle S_{z}\rangle\sin\theta-\langle S_{x}\rangle\cos\theta\right)$
$\hbar\dot{\rho}_{0}$ $=2\mbox{Re}(\Gamma_{g0,0g}^{P_{\alpha}=0}(t))(1-\rho_{0}-\rho_{2})+4\sum_{\alpha}\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\langle S_{z}\rangle/\hbar-4\mbox{Re}(\Gamma_{g0,0g}^{P_{\alpha}=0}(t))\rho_{0}$
$\hbar\dot{\rho}_{2}$ $=2\mbox{Re}(\Gamma_{g2,2g}^{P_{\alpha}=0}(t))(1-\rho_{0}-\rho_{2})-4\sum_{\alpha}\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\langle S_{z}\rangle/\hbar-4\mbox{Re}(\Gamma_{g2,2g}^{P_{\alpha}=0}(t))\rho_{2}.$
The notation $\Gamma_{lj,uv}^{P_{\alpha}=0}(t)$ indicates zero spin polarization for the rates in Eq. 6. It is important to emphasize that the rates $\Gamma_{ev,vg}(t)$, that “flip” the excited state to the ground one or vice versa, are proportional to the spin polarization and $\sin\theta$, making them a consequence of both the lead spin polarization and the homodyne angle, i.e. those rates are exactly zero when $P_{L}=0$ and $\theta=0$. Thus $2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta$ does not depend on $\theta$.
Equations LABEL:spin_equations introduces the spin relaxation rate $1/\tau_{\mathrm{rel}}$ and the non-equilibrium spin accumulation $\langle\dot{S}_{z}\rangle_{\mathrm{acc}}$, parallel to the spin polarization as defined in Ref. [38],
@@
$\langle\dot{S}_{z}\rangle_{\mathrm{acc}}$ $=\sum_{\alpha}P_{\alpha}\Bigg{[}2\mbox{Re}(\Gamma_{0g,g0,\alpha}^{P_{\alpha}=0}(t))\rho_{0}-2\mbox{Re}(\Gamma_{2g,g2,\alpha}^{P_{\alpha}=0}(t))\rho_{2}$ (19)
$+(\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t)-\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t)))(1-\rho_{0}-\rho_{2})\Bigg{]},$
The spin accumulation term in Eq. 19 depends on the populations of transient and single occupied states of the QI, and describes the spin accumulation via tunneling to and from spin-polarized leads, which renders it proportional to the current. In short, the spin accumulation term is a source term that builds up the spin polarization on the QI. The spin relaxation rate Eq. 18 describes the decay of every spin component towards the steady state by electrons with a specific spin tunneling out of the QI, a process with $v=0$, or by tunneling in of a second electron with an opposite spin state, forming a doubly occupied state ($v=2$). The spin accumulation and relaxation govern the spin dynamics of the QI if the spin polarization and magnetic field are parallel, $\theta=0$. However, this is not the complete picture of the dynamics since the angle $\theta$ enters Eqs. LABEL:spin_equations. This orientation dependence can be encompassed in the components of the following rotational term:
$\langle\mathbf{\dot{S}}\rangle_{\mathrm{rot}}$ $=\mathbf{S}\times\sum_{\alpha}g\mu_{B}\mathbf{B_{\alpha}}(t)/\hbar$ (20)
A direct comparison with Eq. LABEL:spin_equations shows that:
@@
$g\mu_{B}\mathbf{B_{exch}}(t)={2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta}\ \mathbf{z}$ (22)

### v2

@@ II.2 Quantum master equation (QME)
$\Gamma_{vl,ju,\alpha}^{-}(t)$ $=$ $\sum_{\sigma}\lambda_{lv,\sigma}^{*}\lambda_{ju,\sigma}\gamma_{\alpha\sigma}(t){\cal I}_{\alpha}^{-}(\Delta_{ju}).$ (8)
Equations (7) and (8) contain the coupling $\gamma_{\alpha\sigma}(t)=\left(1+A_{\alpha}\cos(\omega t)\right)^{2}\gamma_{\alpha\sigma}^{0}$ where $\gamma_{\alpha\sigma}^{0}=2\pi|t^{0}_{\alpha}|^{2}D_{\alpha}(1/2+\sigma P_{\alpha})=\gamma_{\alpha}^{0}(1/2+\sigma P_{\alpha})$ represents the level broadening due to the time-independent hopping, $t^{0}_{\alpha}$, for spin $\sigma$ in the electrode $\alpha$. The Fermi distribution function $f_{\alpha}(\epsilon)=1/\left(e^{\beta_{\alpha}(\epsilon-\mu_{\alpha})}+1\right)$, being $\beta_{\alpha}=1/k_{B}T_{\alpha}$ and $\mu_{\alpha}$ the electrochemical potential, is encoded in the integrals ${\cal I}^{\pm}_{\alpha}(\Delta_{ju})$ which are of the form:
${\cal I}_{\alpha}^{\pm}(\Delta_{ju})$ $=$ $\pm\frac{i}{2\pi}\frac{1}{1+A_{\alpha}\cos(\omega t)}\int^{\infty}_{-\infty}d\epsilon f^{\pm}_{\alpha}(\epsilon)$ (9)
@@
Notice that if $\hbar\omega\ll|\Delta_{ju}|$, Eq. (9) simplifies to [^3] A feasible assumption since $\hbar\omega$ will be in the 0.1 meV range or below, while $\Delta_{ju}\sim 100$ meV. This is because, in the rates only, $\Delta_{ju}$ always represents an energy difference between charged states with different numbers of electrons; otherwise, the rate is zero due to the definition of $\lambda_{lj,\sigma}$. Conversely, $\Delta_{lj}$ in Eq. (5) can be in the $\hbar\omega$ range.:
${\cal I}_{\alpha}^{\pm}(\Delta_{ju})=\pm\frac{i}{2\pi}\int^{\infty}_{-\infty}d\epsilon f^{\pm}_{\alpha}(\epsilon)\frac{1}{\epsilon\mp\Delta_{ju}\pm i\hbar/\tau_{c}},$ (10)
approximation valid for all the simulations in this work. Equation (10) contains $f^{+}_{\alpha}(\epsilon)=f_{\alpha}(\epsilon)$ and $f^{-}_{\alpha}(\epsilon)=1-f_{\alpha}(\epsilon)$ following [41]. Further details on the transition rates, the derivation of the quantum master equation (QME), and its solution in the long-time limit can be found in [37], albeit with some differences, as the theory and physical interpretations have evolved over time.
It is important to mention that $\hbar/\tau_{c}$, up to first order in perturbation theory, should be infinitesimal in Eqs. (9) and (10). However, in this work, $\hbar/\tau_{c}\ll|\Delta_{ju}|\approx|\varepsilon|\ \mathrm{or}\ \varepsilon+U$ but nonzero, allowing us to explore two distinct regimes for a first-order perturbation theory: the one above the thresholds $|\varepsilon|,\varepsilon+U$ where an full electron can hop in or out from the QI and the one below the thresholds where the QI is in the so called Coulomb blockade (CB) regime and we tunnel through the tail of the resonance [38]. Therefore, the transition between both regimes is controlled by the DC voltages and happens when it equalizes the ionization energy $\varepsilon$ or the charging one $\varepsilon+U$, which correspond to the situation of removing and adding one electron on the QI respectively.
We are aware of the limitations a finite $\hbar/\tau_{c}$ introduces, as a proper characterization of the CB requires the use of cotunneling transport or a Kondo scattering Hamiltonian [45, 32]. However, as long as $\hbar/\tau_{c}\ll|\varepsilon|,\varepsilon+U$, we can overcome this limitation of the model and explore, to some extent, the implications of the FLT regime that governs the ESR signal in the CB regime. In this context, we expect low current at low DC biases, consistent with previous studies [41, 70, 45, 71, 46]. Future extensions of this model will aim to circumvent this issue for a better characterization of the system. However, first-order perturbation theory already provides rich results that follow experimental evidence, as we will explore next.
@@ II.3 Expression for the current
This expression corresponds to the lowest order in the coupling (sequential tunneling) and is therefore expected to decrease rapidly in the CB regime, while remaining large and constant outside of it. Equation (11) contains both DC and AC terms as a consequence of the hopping modulation. The measurable quantity in the experiment is the electric current signal in the long-time limit, which depends on the resonance condition as well as the relative orientation of the QI spin to the direction of the tip spin polarization. We refer to this current as ESR signal in the following. The coherences (off-diagonal elements of the density matrix) are explicitly included in the expression and are crucial for the correct calculation of the ESR signal, as pointed out in a previous work [38]. The coherences in Eq. (11) also lead to the so called homodyne detection of the electric current, a product of the misalignment of the tip’s spin and the spin of the impurity under it [42, 22, 73]. Following Appendix C one arrives to:
$I_{L}(t)$ $=$ $\frac{e}{\hbar}\gamma_{L}\left(1+A_{L}\cos(\omega t)\right)^{2}$ (12)
$\times$ $\left[(\rho_{1}-2P_{L}\langle s_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{+}(\Delta_{2g})\right\}\right$.
- $\left.(\rho_{1}+2P_{L}\langle s_{z}\rangle)\mbox{Re}\left\{{\cal I}_{L}^{-}(\Delta_{0g})\right\}\right]$
$=$ $-I_{R}(t)=I(t).$
@@
$I_{L}=-I_{R}$ is due to electric current conservation and $\rho_{1}=\rho_{gg}+\rho_{ee}\approx 1$ if the populations of the transient states ($0,2$) are negligible, which we assume in the following. Equation (12), employs again the notation $g$ to refer to the ground state for a more general treatment instead of $\downarrow$, accounting for the situation in which the spin is tilted by the magnetic field with respect to the spin polarization on $z$. We do this because the eigenstates are always defined in the axis where the QI Hamiltonian is diagonal, see Appendix A, parallel to the magnetic field here, as we have mentioned before. As a consequence of choosing the $z$-axis as the magnetization direction of the leads, where the spin polarization lies, the electric current becomes a measurement of the expectation value $\langle s_{z}\rangle$ as Eq. (12) shows. In general, for any magnetization orientation of the spin-polarized lead, we expect that the current will measure the spin component parallel to it, in agreement with the physical interpretation in the experiments  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30].
@@
Assuming that the magnetic field lies in the $xz$ plane, $\mathbf{B}=B(\sin\theta,0,\cos\theta)$, see Appendix A, we obtain:
$\langle s_{z}\rangle$ $=$ $\frac{\hbar}{2}(\rho_{ee}-\rho_{gg})\cos\theta+\frac{\hbar}{2}(\rho_{ge}+\rho_{eg})\sin\theta,$ (13)
@@
which does not depend on $\rho_{0}$ and $\rho_{2}$, the populations of the unoccupied and doubly states, since they provide zero spin to the simplest QI. $\theta$, also referred to as homodyne angle, is the angle formed between the direction of the left electrode’s magnetization, fixed along the $z$-axis, and the impurity spin when the $g$-factor is isotropic. We should emphasize that in Eq. (13), the main contribution to the populations comes from the Floquet number $n=0$, while the contribution from the coherences arises from $n=1$ [38]. Thus, the coherences are intrinsically time-dependent in the long-time limit. For the sake of completeness, the in-plane components of the spin expectation value are:
@@
The reader might wonder why the expectation value of the spin along $z$ depends on the coherences and not only on the difference in populations or why the $x$-component depends on the population. The reason is the inclusion of mixed states as a consequence of the magnetic field being tilted from the $z$ axis. Mathematically, Eq. (13) arises from changing the basis set from the classical up and down states of the Pauli matrix to the eigenstate basis set of the quantum impurity (QI) since ${\hat{s}}_{i}$ has to be expressed in the same basis as the reduced density matrix, meaning that we cannot simply use the standard Pauli matrices representation. This can be interpreted as a projection of the populations and coherences, defined in the axis where the QI Hamiltonian is diagonal, onto the Cartesian coordinates. Equivalently, one could rotate the cartesian coordinates to the magnetic field direction, changing the basis of the reduced density matrix instead, a procedure done in the literature [42, 22, 26, 19, 74]. We emphasize that $\langle\mathbf{s}\rangle/\hbar$ describes a quantum-statistical average and every component can adopt any number between 1/2-1/2 and $1/2$.
Introducing Eq. (13) into Eq. (12) allows us to obtain an electric current equation that clearly shows the dependence on $\theta$. To simplify the result, we place ourselves in the long time where the populations are constant, i.e. $\rho_{ee(gg)}=P_{e(g)}$. Additionally, we consider that the linear driving term dominates ($A_{L}^{2}\sim 0$) so that the coherences are continuously Larmor precessing only, i.e. $\rho_{ge}(t)=\rho_{ge,-1}e^{i\omega t}+\rho_{ge,1}e^{-i\omega t}$, where the $\pm 1$ in the coherences refers to its Floquet number [38]. This yields:
@@
- $P_{L}\left[(P_{e}-P_{g})\cos\theta+2A_{L}(\rho_{ge,1}+\rho_{eg,-1})\sin\theta\right]$
@@
$I_{L}$ $\approx$ $I_{0}-\underbrace{P_{L}\left[\Delta P_{eg}\cos\theta+2A_{L}(\rho_{ge,1}+\rho_{eg,-1})\sin\theta\right]I_{p}}_{\Delta I}$
where the new definitions $I_{0},\ I_{p}$, containing a non-trivial DC bias dependence, see Eqs. (9) and (10), and $\Delta P_{eg}$ can be deduced by comparing terms. Later on, we conveniently write $\Delta I=I_{L}-I_{\mathrm{BG}}$, where the background current (i.e., baseline or off-resonant current) is defined as $I_{\mathrm{BG}}=I_{L}-\Delta I=I_{0}-2\Delta I$, which depends on the spin polarization $P_{L}$ and the homodyne angle $\theta$.
Equation (LABEL:homo_current_final) highlights the role of spin polarization. For $P_{L}=0$, the electric current simply has a constant unpolarized value, originating from electron or hole hopping from the left to right (or vice versa) through the QI. In this case, the current is independent of the QI populations and coherences, rendering it insensitive to changes in the QI state. Consequently, it cannot induce spin flips or resonant behavior, even under AC driving. A finite spin polarization $P_{L}\neq 0$ breaks this situation, producing an imbalance between up and down states that modifies the tunneling magnetoresistance.
In this regime, a polarized electric current emerges and $\langle s_{z}\rangle$ becomes relevant in Eq. (LABEL:homo_current_final). On resonance and in the long time limit in the CB regime, the two singly occupied levels are equally populated ($1/2$ each) and $\langle s_{z}\rangle\rightarrow 0$, assuming finite but small coherences. By contrast, for $P_{L}\neq 0$ and $A_{L}=0$, the QI retains a net spin polarization $\langle s_{z}\rangle\neq 0$, but the absence of coherences prevents population exchange, and no resonance occurs at the Larmor frequency. Equation (LABEL:homo_current_final) also reveals the dependence on the homodyne angle $\theta$. For $\theta=0$, $\sin\theta=0$, no population exchange takes place, mirroring the undriven case. At $\theta=\pi/2$, the current becomes independent of the populations and the ESR signal is governed solely by the coherences. Thus, Eq. (LABEL:homo_current_final) demonstrates that homodyne detection arises from both the misalignment between the tip and impurity spin and the presence of AC response being measured into the DC current.
@@ II.4 Mapping onto a spin problem
Following Ref. [41], we now aim to map the QME, Eq. (5), onto a system of equations for the spin. This will become particularly useful for building some physical intuition on the definitions of the Rabi rate and the renormalization of the resonance energy. Using Eqs. (13), (14) and (15), the system of equations for the spin components and the population of the transient states is [^4] Equations (17a) and (17b) assume a negligible Zeeman energy in the rates compared to the ionization and charge energy, $\varepsilon$ and $\varepsilon+U$ in order to simplify the equations. This is valid since these energy thresholds are $\sim 100$ meV or more while $\Delta_{eg}=g\mu_{B}B<0.1$ meV in the STM-ESR context  [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. This allows us to write $\sum_{\alpha}P_{\alpha}\left[\mbox{Im}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}-\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0})\right]\sin\theta=\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))$, which further simplifies the spin system of equations; see Appendix EHowever, in the numerical simulations, the Zeeman energy is not neglected anywhere. :
$\hbar\langle\dot{s}_{z}\rangle$ $=\Delta_{ge}\langle{s}_{y}\rangle\sin\theta-\hbar\langle{s}_{z}\rangle/\tau_{\mathrm{rel}}+\hbar\langle\dot{s}_{z}\rangle_{\mathrm{acc}}$ (17a)
$\hbar\langle\dot{s}_{x}\rangle$ $=\left[2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta-\Delta_{ge}\cos\theta\right]\langle{s}_{y}\rangle$
$-\hbar\langle{s}_{x}\rangle/\tau_{\mathrm{rel}}$
$\hbar\langle\dot{s}_{y}\rangle$ $=-2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))\langle s_{x}\rangle/\sin\theta-\hbar\langle{s}_{y}\rangle/\tau_{\mathrm{rel}}$
$-\Delta_{ge}\left(\langle s_{z}\rangle\sin\theta-\langle s_{x}\rangle\cos\theta\right)$
$\hbar\dot{\rho}_{0}$ $=2\mbox{Re}(\Gamma_{g0,0g}^{P_{\alpha}=0}(t))(1-\rho_{0}-\rho_{2})-4\mbox{Re}(\Gamma_{g0,0g}^{P_{\alpha}=0}(t))\rho_{0}$ (17b)
$+4\sum_{\alpha}\mbox{Re}(\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\langle s_{z}\rangle/\hbar$
$\hbar\dot{\rho}_{2}$ $=2\mbox{Re}(\Gamma_{g2,2g}^{P_{\alpha}=0}(t))(1-\rho_{0}-\rho_{2})-4\mbox{Re}(\Gamma_{g2,2g}^{P_{\alpha}=0}(t))\rho_{2}$
$-4\sum_{\alpha}\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t))P_{\alpha}\langle s_{z}\rangle/\hbar.$
The notation $\Gamma_{lj,uv}^{P_{\alpha}=0}(t)$ indicates zero spin polarization for the rates in Eq. (6). It is important to emphasize that the rates $\Gamma_{ev,vg}(t)$, that “flip” the excited state to the ground one or vice versa, are proportional to the spin polarization and $\sin\theta$, making them a consequence of both the lead spin polarization and the homodyne angle, i.e. those rates are exactly zero when $P_{L}=0$ and $\theta=0$. Thus $2\sum_{v=0,2}\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta$ does not depend on $\theta$.
Equations (17a) introduces the spin relaxation rate $1/\tau_{\mathrm{rel}}$ and the non-equilibrium spin accumulation $\langle\dot{s}_{z}\rangle_{\mathrm{acc}}$, parallel to the spin polarization as defined in Ref. [41],
@@
$\langle\dot{s}_{z}\rangle_{\mathrm{acc}}$ $=2\sum_{\alpha}P_{\alpha}\Bigg[\mbox{Re}(\Gamma_{0g,g0,\alpha}^{P_{\alpha}=0}(t))\rho_{0}-\mbox{Re}(\Gamma_{2g,g2,\alpha}^{P_{\alpha}=0}(t))\rho_{2}$ (19)
$+(\mbox{Re}(\Gamma_{g2,2g,\alpha}^{P_{\alpha}=0}(t)-\Gamma_{g0,0g,\alpha}^{P_{\alpha}=0}(t)))(1-\rho_{0}-\rho_{2})/2\Bigg],$
The non-equilibrium spin accumulation term, Eq. (19), depends on the populations of transient and single-occupied states of the QI, and describes the spin accumulation via tunneling to and from spin-polarized leads, which renders it proportional to the current. In short, the spin accumulation term is a source term that builds up the spin polarization on the QI. The time-independent spin accumulation defines a characteristic timescale, $\hbar/\langle\dot{s}_{z}\rangle_{\mathrm{acc}}$, over which it acts on the spin. For DC voltages within the CB regime, the effect of the spin accumulation can be partly neglected, since spin relaxation time is much shorter than the spin accumulation time and dominates the spin dynamics.
The spin relaxation rate Eq. (18) describes the decay of every spin component towards the steady state by electrons with a specific spin tunneling out of the QI, a process with $v=0$, or by tunneling in of a second electron with an opposite spin state, forming a doubly occupied state ($v=2$). The spin accumulation and relaxation govern the spin dynamics of the QI if the spin polarization and magnetic field are parallel, $\theta=0$. However, this is not the complete picture of the dynamics since the angle $\theta$ enters Eqs. (17a). This orientation dependence can be encompassed in the components of the following rotational term:
$\langle\mathbf{\dot{s}}\rangle_{\mathrm{rot}}$ $=\langle\mathbf{s}\rangle\times\sum_{\alpha}g\mu_{B}\mathbf{B_{\alpha}}(t)/\hbar$ (20)
A direct comparison with Eq. (17a) shows that:
@@
$g\mu_{B}\mathbf{B_{exch}}(t)=2\sum_{v=0,2}\left[\mbox{Im}(\Gamma_{ev,vg}(t))/\sin\theta\right]\mathbf{z}$ (22)